OpenStudy (anonymous):

Solve the triangle. B = 36°, a = 38, c = 17

OpenStudy (jdoe0001):

hmmm what "solve" means?

OpenStudy (anonymous):

find b, A, and C I have found "b" b=26.2

OpenStudy (anonymous):

the answer options are: b ≈ 41.3, C ≈ 22.3, A ≈ 121.7 b ≈ 41.3, C ≈ 26.3, A ≈ 117.7 b ≈ 26.2, C ≈ 118.7, A ≈ 25.3 b ≈ 26.2, C ≈ 22.3, A ≈ 121.7

OpenStudy (jdoe0001):

I assume you've covered the Law of Cosines?

OpenStudy (anonymous):

and then I used the Law of Cosine to find "A," but I get =189.40 which isn't a choice

OpenStudy (jdoe0001):

hmm what did you get for side "b"?

OpenStudy (anonymous):

For side "b" I got 26.2... by using the equatoin b^2 = a^2 + c^2 - 2ac(CosB)

OpenStudy (jdoe0001):

ok I got 26.2 as well

OpenStudy (anonymous):



OpenStudy (anonymous):

now use sin formula to find the angle A or angle C or both.

OpenStudy (jdoe0001):

\(\bf \textit{Law of Cosines}\\ \quad \\ a^2 = {\color{blue}{ b}}^2+{\color{red}{ c}}^2-(2{\color{blue}{ b}}{\color{red}{ c}})cos(A)\implies cos(A)=\cfrac{{\color{blue}{ b}}^2+{\color{red}{ c}}^2-a^2}{2{\color{blue}{ b}}{\color{red}{ c}}}\\ \quad \\ \measuredangle A = cos^{-1}\left(\cfrac{{\color{blue}{ b}}^2+{\color{red}{ c}}^2-a^2}{2{\color{blue}{ b}}{\color{red}{ c}}}\right)\)

OpenStudy (anonymous):

\[\frac{ 26.2 }{ \sin 36 }=\frac{ 38 }{ \sin A }=\frac{ 17 }{ \sin C }\]

OpenStudy (anonymous):

@surjithayer and @jdoe0001 Thank you both so much! This is my first time on here and I didn't think everyone would be so nice! :)

OpenStudy (anonymous):

yw